How do i find the Market value of debenture

R1 200 000 8% debentures due in 6 years’ time and the current yield-to-maturity is 6%
Answer #1

Calculation of WACC by using Growth model for calculation of ke-

ke= DPS1/P0 +g

DPS1= DPS0(1+g)

=0.30(1+.13)

=.339

Ke = (.339/4.5) +.13

=20.5%

Kp= 14%

Kd=Interest rate(1-tax)

=8(1-.3)

=5.6%

Kl= 13(1-.3)

=9.1%

CapitalMVWeightcostw×c
Equity9000000.48620.59.963
Reserves5000000.27020.55.535
Debt1200000.0655.6.364
Preference share capital2500000.135141.89
Bank loan800000.0439.1.391

WACC= 18.143%

CALCULATION OF WACC IF KE IS CALCULATED USING CAPM APPROACH-

ke= rf+ beta(market risk premium)

= 5+1.7(8)

=18.6%

Kd, kl, kp will remain same.

Capital weight cost w×c

Equity .486 18.6 9.04

WACC under such =18.143(from above) -9.963(calculated using growth model -eliminated)

+9.04(required)

=17.22%


answered by: ANURANJAN SARSAM